What concentration of glucose is isosmotic to a 0.25 M solution of NaCl?
1
Expert's answer
2017-02-21T07:49:13-0500
Osmolarity of NaCl = i*C*R*T, where i=2 because NaCl disosiated into two ions.
Osmolarity of glucose = i*C*R*T, where i=1 brcause glucose does not dissosiate.

Thus, concentration of glucose = 2*0.25*R*T/1*R*T = 0.5 M